# 00footnotetext: Received xxxx June xxxx

### v1

@@ 8 Results
As no significant signal is observed in data, the upper limit on $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})$ is set using a Bayesian method described in Ref. [58]. We perform a series of maximum likelihood fits as described in Section 5, with the signal yield $N_{\rm DT}$ fixed to each scanned value. The corresponding maximum likelihood values are used to construct a discrete likelihood distribution, $\mathcal{L}(\mathcal{B}_{\rm sig})$. The systematic uncertainties are incorporated in two steps: first, the additive uncertainties are addressed by varying the DT fit method, with the most conservative upper limit retained. Next, the multiplicative uncertainties are accounted for by smearing the likelihood distribution as [59]
@@
where $\epsilon_{0}$ is the nominal signal efficiency and $\sigma_{\epsilon}$ is the multiplicative uncertainty corresponding to the efficiency value. By integrating the $\mathcal{L}^{\prime}(\mathcal{B}_{\rm sig})$ curve up to 90% of the area for $N_{\rm DT}>0$ and calculating the corresponding BF using Eq. (2), we set the upper limit on the BF of $D^{+}\to\gamma e^{+}\nu_{e}$ at 90% C.L. to be $1.2\times 10^{-5}$, as shown in Figure 8,
@@ 9 Summary
In conclusion, we have conducted an improved search for the radiative leptonic decay $D^{+}\to\gamma e^{+}\nu_{e}$ using 20.3 $\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ collision data taken at $\sqrt{s}=3.773$ GeV with the BESIII detector. Our investigation does not reveal a significant signal, and establish an upper limit on its BF with radiative photon energy over 10 MeV to be $1.2\times 10^{-5}$ at 90% C.L.. This result excludes most existing theoretical predictions [16, 17, 18, 19, 10], thereby indicating the necessity of further investigation of the decay mechanism. The key advancement of this study is the application of a novel deep learning approach, which effectively distinguishes the signals from massive backgrounds. Our results showcase the potential and wide-ranging applicability of deep learning techniques in particle physics research.

### v2

@@ 8 Results
Given that no significant signal was observed in data, the upper limit on $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})$ was set using a Bayesian method described in Ref. [58]. We performed a series of maximum likelihood fits as described in Section 5, with the signal yield $N_{\rm DT}$ fixed to each scanned value. The corresponding maximum likelihood values were used to construct a discrete likelihood distribution, $\mathcal{L}(\mathcal{B}_{\rm sig})$. The systematic uncertainties were incorporated in two steps: first, the additive uncertainties were addressed by varying the DT fit method, with the most conservative upper limit retained. Next, the multiplicative uncertainties were considered by smearing the likelihood distribution as follows [59]
@@
where $\epsilon_{0}$ is the nominal signal efficiency, and $\sigma_{\epsilon}$ is the multiplicative uncertainty corresponding to the efficiency value. By integrating the $\mathcal{L}^{\prime}(\mathcal{B}_{\rm sig})$ curve up to 90% of the area for $N_{\rm DT}>0$ and calculating the corresponding BF using Eq. (2), we set the upper limit on the BF of $D^{+}\to\gamma e^{+}\nu_{e}$ at 90% C.L. to be $1.2\times 10^{-5}$, as shown in Figure 8,
@@ 9 Summary
We conducted an improved search for the radiative leptonic decay $D^{+}\to\gamma e^{+}\nu_{e}$ using 20.3 $\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ collision data taken at $\sqrt{s}=3.773$ GeV with the BESIII detector. Our investigation does not reveal a significant signal but establishes an upper limit of $1.2\times 10^{-5}$ on its BF with a radiative photon energy above 10 MeV, at 90% C.L.. This result excludes most existing theoretical predictions [16, 17, 18, 19, 10], thereby indicating the necessity for further investigation of the decay mechanism. The key advancement of this study is the application of a novel deep learning approach, which effectively distinguishes the signals from massive backgrounds. Our results showcase the potential and broad applicability of deep learning techniques in particle physics research.
10 Acknowledgment
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support.